Eagle Company makes the MusicFinder, a sophisticated satellite radio. Eagle has experienced a steady growth in sales for the past five years. However, Ms. Luray, Eagle's CEO, believes that to maintain the company's present growth will require an aggressive advertising campaign next year. To prepare for the campaign, the company's accountant, Mr. Bednarik, has prepared and presented to Ms. Luray the following data for the current year, year 1:

Variable costs:
Direct labor (per unit) $ 81
Direct materials (per unit) 37
Variable overhead (per unit) 12
Total variable costs (per unit) $ 130
Fixed costs (annual):
Manufacturing $ 390,000
Selling 293,000
Administrative 796,000
Total fixed costs (annual) $ 1,479,000
Selling price (per unit) 412
Expected sales revenues, year 1 (28,000 units) $ 11,536,000

Eagle has an income tax rate of 35 percent.

Ms. Luray has set the sales target for year 2 at a level of $13,184,000 (or 32,000 radios).

Required:

a. What is the projected after-tax operating profit for year 1?

b. What is the break-even point in units for year 1? (Round up your answer to the nearest whole number.)


c. Ms. Luray believes that to attain the sales target (32,000 radios) will require additional selling expenses of $294,000 for advertising in year 2, with all other costs remaining constant. What will be the after-tax operating profit for year 2 if the firm spends the additional $294,000?

d. What will be the break-even point in sales dollars for year 2 if the firm spends the additional $294,000 for advertising? (Solve by computing volume in units first. Round up units to the nearest whole number and round your final answer to the nearest whole dollar amount.)

e. If the firm spends the additional $294,000 for advertising in year 2, what is the sales level in dollars required to equal the year 1 after-tax operating profit? (Solve by computing volume in units first. Round up units to the nearest whole number and round your final answer to the nearest whole dollar amount.)
Answer #1

a projected after tax operating profit year I particulars Amount (in & $ 11,536,000 Sales revenue - less: Variable cost totalb) Break even point in units for year ! Break even in units = Total Fixed cost contribution per unit 1,479,000 (7896000/28000@ year-2 Operating profit after tax Particulars Amant (in $) Sales [32000x412) -> $13,184,000 lesso Variable cost - > $ (4.16Break even sales 1723,000 1 Break even sales in units = (9024000/2000 = 6287.23 =6287 units Break even sales = 62874412 +3599